# Can Current Detectors Catch Face-to-Voice Deepfake Attacks?

### v1

@@ 1 Introduction
Recent advances in generative AI have enabled speech synthesis models capable of mimicking human voices with remarkable realism. While these systems have legitimate uses in entertainment and accessibility, they also introduce serious risks to security, privacy, and trust. Synthetic voices are increasingly used to impersonate individuals, bypass authentication systems, and facilitate fraud[14]. To counter these risks, researchers have developed a growing body of audio deepfake detectors [2, 12, 17]. However, existing research on audio deepfakes has largely focused on text- or voice-conditioned pipelines, where artifacts left by vocoders provide detectable cues [22, 23]. The emergence of new cross-modal synthesis methods has begun to erode these assumptions.
A particularly concerning example is Face-to-Voice (FOICE) synthesis [9], which generates a synthetic voice from a single facial image by exploiting correlations between facial and vocal characteristics, as illustrated in Figure 1. This enables large-scale impersonation attacks without needing any clean voice samples, since facial images are readily available on social media. For example, an attacker could scrape a victim’s photo and phone number (often used as a WeChat ID) from social media such as LinkedIn and, using an SV2TTS‑style vocoder, render attacker‑supplied text (e.g., authentication digits) in a voice closely resembling the target, producing speech convincing enough to bypass authentication systems like WeChat Voiceprint and Microsoft Azure. These capabilities fundamentally shift the threat landscape: attackers are no longer constrained by voice data availability, while existing detectors remain optimized for vocoder-driven synthesis.
@@ 2.1 Audio Deepfakes and Detection Landscape
The problem of detecting synthetic speech has been studied extensively in recent years. Early work focused on handcrafted spectral, phase, and prosodic features combined with machine learning classifiers, formalized through shared tasks such as the ASVspoof challenges [20]. These benchmarks remain central for evaluation, but detectors trained solely on them often inherit dataset-specific biases. Subsequent corpora, such as WaveFake [7], LibriSeVoc [15], and FakeAVCeleb [10], broadened vocoder and synthesis diversity, yet cross-domain generalization remains limited.
Modern detectors increasingly adopt deep neural architectures. Convolutional and recurrent models applied to spectrograms, as well as waveform-level networks, dominate the landscape. More recent designs leverage attention or graph-based modeling to capture both local and global structure, with AASIST [21] as a representative state-of-the-art detector. Parallel efforts explicitly target artifacts introduced during speech synthesis, e.g., residual noise or vocoder frequency patterns [18], while others explore domain-invariant representations [15] and pretraining with large-scale multimodal encoders [5]. Temporal–channel modeling [19] has also been investigated to improve robustness against noisy playback and unseen vocoder types.
Despite these advances, most detectors remain tuned to the artifact patterns of vocoder-based text-to-speech or voice-conversion pipelines, limiting their ability to generalize to alternative generation mechanisms. As a result, detectors degrade sharply when confronted with unseen generation methods, such as SpeechT5 [1], or under realistic environmental noise conditions. Crucially, no existing benchmark or study has systematically evaluated detection against face-conditioned, cross-modal synthesis such as FOICE.
@@ 2.2 FOICE and Face-to-Voice Synthesis
FOICE, introduced at USENIX Security 2024 [9], exemplifies a qualitatively new and stealthier threat model. Instead of relying on text or voice conditioning, FOICE infers how a person would sound directly from a single facial image. As shown in Figure 1, it exploits the biological correlation between facial structure (e.g., jawline, lips, bone morphology) and vocal attributes (e.g., pitch, timbre, resonance). The model separates face-dependent voice features from face-independent features sampled from a latent Gaussian space. These are fused and synthesized through a SV2TTS-style vocoder, which renders attacker-provided text in the inferred voice. This design makes FOICE stealthy and scalable: facial images are far easier to obtain than voice recordings, and FOICE has been shown to succeed even from low-resolution or partially occluded faces [9]. Because FOICE is conditioned on facial rather than vocal inputs, its outputs lack many of the frequency-domain artifacts that detectors rely on, creating a serious blind spot for current defenses. In evaluations, FOICE achieved average per-victim bypass success rates of $\sim$30% against WeChat Voiceprint and nearly 100% overall success when multiple attempts were allowed [9].
@@ 3.2 Detection Models
(1) AASIST [21]. A waveform-level model that combines spectro–temporal front-ends with graph attention to capture both local and global dependencies. Its architecture makes it a strong candidate for detecting cross-modal synthesis artifacts beyond traditional vocoder traces.
(2) Ren et al. [15]. A model designed to separate domain-specific from domain-invariant features to support cross-domain generalization. We expect this model to be informative for testing resilience to unseen pipelines (e.g., SpeechT5) and to resist catastrophic forgetting under FOICE fine-tuning.
(3) Sun et al. [18]. A spectrogram-based model trained to identify residual cues left by neural vocoders. Although effective on vocoder-based synthesis, it serves here as a stress-test for FOICE, which lacks many such artifacts.
(4) Temporal–Channel Modeling (TCM) [19]. A self-attention-based model that captures by attending jointly over temporal and channel dimensions. It provides a strong attention-based baseline under both in-distribution and out-of-distribution conditions.
@@ 3.3 Datasets
(A) VoxCeleb2 [3]. Contains over one million utterances from thousands of celebrity speakers recorded in diverse recording environments. Used both to generate FOICE voices and to evaluate robustness on in-distribution real speech.
(B) AVSpeech [6]. Comprises large-scale audio-visual speech segments from YouTube, with cleaner acoustic conditions and more diverse (non-celebrity) speaker identities. Enables evaluation of generalization beyond FOICE’s celebrity-centric training distribution.
@@
(D) SpeechT5 DS [1]. Includes 1,400 synthetic samples generated using SpeechT5 (which leverages HiFi-GAN [11]). Used exclusively at evaluation time to test cross-domain generalization (RQ2); it is excluded from baseline training and FOICE fine-tuning stages.
Overall, the dataset comprises 15,144 samples: 6,261 labeled as fake and 8,883 as real. All real speech comes from VoxCeleb2 and AVSpeech, while synthetic samples are split between FOICE and SpeechT5. The dataset is balanced across three playback conditions: clean, noisy (via audiomentations [8]), and denoised (via noisereduce [16]). Real and fake distributions are matched across noise settings to ensure robust evaluation.
@@ 3.4 Fine-Tuning Mechanism
(1) AASIST [21]: batch size 24, 10 epochs, Adam with $\beta=(0.9,0.999)$, CCE loss, base LR $1\times 10^{-4}$, weight decay $1\times 10^{-4}$.
(2) Ren et al. [15]: batch size 8, 30 epochs, cosine annealing, initial LR per original paper, weight decay $1\times 10^{-4}$.
(3) Sun et al. [18]: batch size 32, 5 epochs, LR $1\times 10^{-4}$, weight decay $1\times 10^{-4}$.
(4) TCM [19]: batch size 16, 30 epochs, LR $1\times 10^{-6}$, weight decay $1\times 10^{-4}$.
@@ 4 Evaluation
The vocoder artifact detector by Sun et al. [18] and TCM are the weakest performers: Sun et al. [18] underperforms due to reliance on vocoder residue cues that FOICE does not expose reliably, while TCM exhibits systematic over-detection, resulting in inflated recall at the expense of accuracy, precision, F1-Score across clean, noisy, and denoised conditions. Ren et al.’s model [15] offers the most stable cross-condition performance and achieves the highest F1-score, suggesting strong alignment with the signal characteristics of FOICE speech. AASIST, while slightly behind in F1, maintains a solid balance of sensitivity and specificity, outperforming artifact-centric models on robustness across conditions.
@@
AASIST [21] Base 0.163 0.735 0.416 1.000 0.587 687 1996 965 0
AASIST [21] Fine-tuned 0.003 0.995 0.972 1.000 0.986 687 2941 20 0
Ren et al. [15] Base 0.271 0.898 1.000 0.457 0.627 314 2961 0 373
Ren et al. [15] Fine-tuned 0.002 0.998 0.987 1.000 0.993 687 2952 9 0
Sun et al. [18] Base 0.566 0.630 0.099 0.119 0.108 82 2213 745 605
Sun et al. [18] Fine-tuned 0.004 0.995 0.974 0.999 0.986 686 2940 18 1
TCM [19] Base 0.433 0.297 0.211 1.000 0.349 687 395 2563 0
TCM [19] Fine-tuned 0.150 0.913 0.779 0.750 0.764 515 2812 146 172
@@
RQ2 — Fine-tuning vs. Overfitting. Fine-tuning on FOICE data yields substantial performance gains across all detectors. Previously underperforming models such as TCM and Sun et al. [18] achieve near-perfect accuracy after fine-tuning, as seen in Table I (Model = "Fine-tuned"). Even the stronger models AASIST and Ren et al. [15] improve, suggesting that prior failures stemmed from gaps in training data rather than fundamental architectural limitations. However, these gains often come at the expense of generalization. When evaluated on the out-of-distribution SpeechT5 dataset, fine-tuned models experience significant accuracy drops, as shown in Figure  4. AASIST maintains a stable accuracy while TCM drops by nearly 10%. Sun et al. [18]’s detector nearly collapses (dropping from 38.7% to 3.9% accuracy). In contrast, only Ren et al. [15]’s domain invariant model improves, demonstrating the advantage of architectures explicitly designed for balancing adaptation and generalization.
@@
Key Takeaways. Our findings underscore two critical points. First, state-of-the-art detectors exhibit clear blind spots when confronted with FOICE deepfakes, which fall outside the assumptions of vocoder-based synthesis pipelines. Second, while fine-tuning can yield near-perfect detection within-distribution, it often erodes robustness against unseen synthesis, unless the model architecture is explicitly optimized for domain robustness, as demonstrated by Ren et al. [15]. Addressing this trade-off is essential for building durable defenses against evolving cross-modal deepfake techniques.
@@ 5 Discussion
This fine-tuning–generalization trade-off mirrors long-standing challenges observed in the visual deepfake detection literature. Numerous studies show that detectors trained on specific generators or datasets rarely generalize to unseen manipulations, often overfitting to superficial artifacts or identity-specific cues. For example, Le et al. (EuroS&P 2025) [13] show that none of sixteen state-of-the-art detectors maintained robust cross-generator performance, several even performed worse than random guessing on novel forgeries. Yang et al. (CVPR 2025) [24] similarly demonstrate that detectors optimized on a single generator exhibit poor transfer to others, and propose multi-generator discrepancy learning to improve robustness. Dong et al. (CVPR 2023) [4] further attribute this fragility to implicit identity leakage, where detectors inadvertently entangle artifact detection with identity cues, limiting cross-domain effectiveness. Collectively, these findings reinforce our results on FOICE: while fine-tuning offers strong in-distribution performance, its specialization often comes at the cost of generalization, a systemic limitation shared across both audio and visual modalities.
Error patterns across detectors. An inspection of confusion matrix counts reveals divergent failure modes: baseline TCM misclassifies nearly all real samples as fake, leading to extremely high false positive rates, while base Ren et al. [15] exhibits the opposite pattern with widespread false negatives. These biases underscore that detectors not only vary in performance but also in the nature of their errors, highlighting the importance of examining decision tendencies beyond aggregate metrics.
Detection under noisy and denoised conditions. Our experiments also highlight the impact of noise and audio preprocessing on detection performance. Interestingly, we observe that denoised audio is often more challenging for detectors than noisy audio. While fine-tuned AASIST and Sun et al. [18] models remain relatively stable across both conditions, detectors like TCM and Ren et al.[15] show substantial degradation post-denoising. TCM’s fine-tuned variant, for example, sees accuracy drop below 90% and EER spike to 0.324 on denoised audio. This may suggest that certain detectors rely on fine-grained noise artifacts or phase distortions that are inadvertently removed by denoising algorithms, masking cues needed for detection. Future work should explore whether more adaptive or perceptually-aligned denoising techniques can retain deepfake-relevant artifacts without suppressing critical detection signals.
@@
To address these gaps, we propose a focused research agenda. First, researchers should expand generator coverage by evaluating FOICE variants integrated with alternative vocoders and video-conditioned generators to measure how vocoder choice and multi-modal conditioning affect both attack success and detector transfer. Second, studies must examine robustness under realistic channels, testing detectors across a matrix of codecs, microphone types, playback devices, and network transport conditions (including compression and packet loss) to better approximate operational environments. Finally, we encourage focused mitigation research on multi-generator / discrepancy training (cf.  [24]), domain-invariant representation learning, class-imbalance–aware fine-tuning, and perceptually-aware denoising that preserves forensic cues. These directions directly address the limitations above and point toward defensible, deployment-oriented research priorities.

### v2

@@ 1 Introduction
Recent advances in generative AI have enabled speech synthesis models capable of mimicking human voices with remarkable realism. While these systems have legitimate uses in entertainment and accessibility, they also introduce serious risks to security, privacy, and trust. Synthetic voices are increasingly used to impersonate individuals, bypass authentication systems, and facilitate fraud[survey]. To counter these risks, researchers have developed a growing body of audio deepfake detectors [blue2022usenix, kumari2025voiceradar, sun2023aisynth]. However, existing research on audio deepfakes has largely focused on text- or voice-conditioned pipelines, where artifacts left by vocoders provide detectable cues [yadav2024fairssd, yan2025voicewukong]. The emergence of new cross-modal synthesis methods has begun to erode these assumptions.
A particularly concerning example is Face-to-Voice (FOICE) synthesis [FOICE], which generates a synthetic voice from a single facial image by exploiting correlations between facial and vocal characteristics, as illustrated in Figure 1. This enables large-scale impersonation attacks without needing any clean voice samples, since facial images are readily available on social media. For example, an attacker could scrape a victim’s photo and phone number (often used as a WeChat ID) from social media such as LinkedIn and, using an SV2TTS‑style vocoder, render attacker‑supplied text (e.g., authentication digits) in a voice closely resembling the target, producing speech convincing enough to bypass authentication systems like WeChat Voiceprint and Microsoft Azure. These capabilities fundamentally shift the threat landscape: attackers are no longer constrained by voice data availability, while existing detectors remain optimized for vocoder-driven synthesis.
@@ 2.1 Audio Deepfakes and Detection Landscape
The problem of detecting synthetic speech has been studied extensively in recent years. Early work focused on handcrafted spectral, phase, and prosodic features combined with machine learning classifiers, formalized through shared tasks such as the ASVspoof challenges [wang2020asvspoof]. These benchmarks remain central for evaluation, but detectors trained solely on them often inherit dataset-specific biases. Subsequent corpora, such as WaveFake [wavefake], LibriSeVoc [AISVGeneralization], and FakeAVCeleb [fakeavceleb], broadened vocoder and synthesis diversity, yet cross-domain generalization remains limited.
Modern detectors increasingly adopt deep neural architectures. Convolutional and recurrent models applied to spectrograms, as well as waveform-level networks, dominate the landscape. More recent designs leverage attention or graph-based modeling to capture both local and global structure, with AASIST [AASIST] as a representative state-of-the-art detector. Parallel efforts explicitly target artifacts introduced during speech synthesis, e.g., residual noise or vocoder frequency patterns [VocoderArtifacts], while others explore domain-invariant representations [AISVGeneralization] and pretraining with large-scale multimodal encoders [llm_deepfake]. Temporal–channel modeling [tcm_add] has also been investigated to improve robustness against noisy playback and unseen vocoder types.
Despite these advances, most detectors remain tuned to the artifact patterns of vocoder-based text-to-speech or voice-conversion pipelines, limiting their ability to generalize to alternative generation mechanisms. As a result, detectors degrade sharply when confronted with unseen generation methods, such as SpeechT5 [speecht5], or under realistic environmental noise conditions. Crucially, no existing benchmark or study has systematically evaluated detection against face-conditioned, cross-modal synthesis such as FOICE.
@@ 2.2 FOICE and Face-to-Voice Synthesis
FOICE, introduced at USENIX Security 2024 [FOICE], exemplifies a qualitatively new and stealthier threat model. Instead of relying on text or voice conditioning, FOICE infers how a person would sound directly from a single facial image. As shown in Figure 1, it exploits the biological correlation between facial structure (e.g., jawline, lips, bone morphology) and vocal attributes (e.g., pitch, timbre, resonance). The model separates face-dependent voice features from face-independent features sampled from a latent Gaussian space. These are fused and synthesized through a SV2TTS-style vocoder, which renders attacker-provided text in the inferred voice. This design makes FOICE stealthy and scalable: facial images are far easier to obtain than voice recordings, and FOICE has been shown to succeed even from low-resolution or partially occluded faces [FOICE]. Because FOICE is conditioned on facial rather than vocal inputs, its outputs lack many of the frequency-domain artifacts that detectors rely on, creating a serious blind spot for current defenses. In evaluations, FOICE achieved average per-victim bypass success rates of $\sim$30% against WeChat Voiceprint and nearly 100% overall success when multiple attempts were allowed [FOICE].
@@ 3.2 Detection Models
(1) AASIST [AASIST]. A waveform-level model that combines spectro–temporal front-ends with graph attention to capture both local and global dependencies. Its architecture makes it a strong candidate for detecting cross-modal synthesis artifacts beyond traditional vocoder traces.
(2) Ren et al. [AISVGeneralization]. A model designed to separate domain-specific from domain-invariant features to support cross-domain generalization. We expect this model to be informative for testing resilience to unseen pipelines (e.g., SpeechT5) and to resist catastrophic forgetting under FOICE fine-tuning.
(3) Sun et al. [VocoderArtifacts]. A spectrogram-based model trained to identify residual cues left by neural vocoders. Although effective on vocoder-based synthesis, it serves here as a stress-test for FOICE, which lacks many such artifacts.
(4) Temporal–Channel Modeling (TCM) [tcm_add]. A self-attention-based model that captures by attending jointly over temporal and channel dimensions. It provides a strong attention-based baseline under both in-distribution and out-of-distribution conditions.
@@ 3.3 Datasets
(A) VoxCeleb2 [chung2018voxceleb2]. Contains over one million utterances from thousands of celebrity speakers recorded in diverse recording environments. Used both to generate FOICE voices and to evaluate robustness on in-distribution real speech.
(B) AVSpeech [ephrat2018looking]. Comprises large-scale audio-visual speech segments from YouTube, with cleaner acoustic conditions and more diverse (non-celebrity) speaker identities. Enables evaluation of generalization beyond FOICE’s celebrity-centric training distribution.
@@
(D) SpeechT5 DS [speecht5]. Includes 1,400 synthetic samples generated using SpeechT5 (which leverages HiFi-GAN [hifi]). Used exclusively at evaluation time to test cross-domain generalization (RQ2); it is excluded from baseline training and FOICE fine-tuning stages.
Overall, the dataset comprises 15,144 samples: 6,261 labeled as fake and 8,883 as real. All real speech comes from VoxCeleb2 and AVSpeech, while synthetic samples are split between FOICE and SpeechT5. The dataset is balanced across three playback conditions: clean, noisy (via audiomentations [audiomentations]), and denoised (via noisereduce [noisereduce]). Real and fake distributions are matched across noise settings to ensure robust evaluation.
@@ 3.4 Fine-Tuning Mechanism
(1) AASIST [AASIST]: batch size 24, 10 epochs, Adam with $\beta=(0.9,0.999)$, CCE loss, base LR $1\times 10^{-4}$, weight decay $1\times 10^{-4}$.
(2) Ren et al. [AISVGeneralization]: batch size 8, 30 epochs, cosine annealing, initial LR per original paper, weight decay $1\times 10^{-4}$.
(3) Sun et al. [VocoderArtifacts]: batch size 32, 5 epochs, LR $1\times 10^{-4}$, weight decay $1\times 10^{-4}$.
(4) TCM [tcm_add]: batch size 16, 30 epochs, LR $1\times 10^{-6}$, weight decay $1\times 10^{-4}$.
@@ 4 Evaluation
The vocoder artifact detector by Sun et al. [VocoderArtifacts] and TCM are the weakest performers: Sun et al. [VocoderArtifacts] underperforms due to reliance on vocoder residue cues that FOICE does not expose reliably, while TCM exhibits systematic over-detection, resulting in inflated recall at the expense of accuracy, precision, F1-Score across clean, noisy, and denoised conditions. Ren et al.’s model [AISVGeneralization] offers the most stable cross-condition performance and achieves the highest F1-score, suggesting strong alignment with the signal characteristics of FOICE speech. AASIST, while slightly behind in F1, maintains a solid balance of sensitivity and specificity, outperforming artifact-centric models on robustness across conditions.
@@
AASIST [AASIST] Base 0.163 0.735 0.416 1.000 0.587 687 1996 965 0
AASIST [AASIST] Fine-tuned 0.003 0.995 0.972 1.000 0.986 687 2941 20 0
Ren et al. [AISVGeneralization] Base 0.271 0.898 1.000 0.457 0.627 314 2961 0 373
Ren et al. [AISVGeneralization] Fine-tuned 0.002 0.998 0.987 1.000 0.993 687 2952 9 0
Sun et al. [VocoderArtifacts] Base 0.566 0.630 0.099 0.119 0.108 82 2213 745 605
Sun et al. [VocoderArtifacts] Fine-tuned 0.004 0.995 0.974 0.999 0.986 686 2940 18 1
TCM [tcm_add] Base 0.433 0.297 0.211 1.000 0.349 687 395 2563 0
TCM [tcm_add] Fine-tuned 0.150 0.913 0.779 0.750 0.764 515 2812 146 172
@@
RQ2 — Fine-tuning vs. Overfitting. Fine-tuning on FOICE data yields substantial performance gains across all detectors. Previously underperforming models such as TCM and Sun et al. [VocoderArtifacts] achieve near-perfect accuracy after fine-tuning, as seen in Table I (Model = "Fine-tuned"). Even the stronger models AASIST and Ren et al. [AISVGeneralization] improve, suggesting that prior failures stemmed from gaps in training data rather than fundamental architectural limitations. However, these gains often come at the expense of generalization. When evaluated on the out-of-distribution SpeechT5 dataset, fine-tuned models experience significant accuracy drops, as shown in Figure  4. AASIST maintains a stable accuracy while TCM drops by nearly 10%. Sun et al. [VocoderArtifacts]’s detector nearly collapses (dropping from 38.7% to 3.9% accuracy). In contrast, only Ren et al. [AISVGeneralization]’s domain invariant model improves, demonstrating the advantage of architectures explicitly designed for balancing adaptation and generalization.
@@
Key Takeaways. Our findings underscore two critical points. First, state-of-the-art detectors exhibit clear blind spots when confronted with FOICE deepfakes, which fall outside the assumptions of vocoder-based synthesis pipelines. Second, while fine-tuning can yield near-perfect detection within-distribution, it often erodes robustness against unseen synthesis, unless the model architecture is explicitly optimized for domain robustness, as demonstrated by Ren et al. [AISVGeneralization]. Addressing this trade-off is essential for building durable defenses against evolving cross-modal deepfake techniques.
@@ 5 Discussion
This fine-tuning–generalization trade-off mirrors long-standing challenges observed in the visual deepfake detection literature. Numerous studies show that detectors trained on specific generators or datasets rarely generalize to unseen manipulations, often overfitting to superficial artifacts or identity-specific cues. For example, Le et al. (EuroS&P 2025) [le2025sok] show that none of sixteen state-of-the-art detectors maintained robust cross-generator performance, several even performed worse than random guessing on novel forgeries. Yang et al. (CVPR 2025) [yang2025d] similarly demonstrate that detectors optimized on a single generator exhibit poor transfer to others, and propose multi-generator discrepancy learning to improve robustness. Dong et al. (CVPR 2023) [Dong_2023_CVPR] further attribute this fragility to implicit identity leakage, where detectors inadvertently entangle artifact detection with identity cues, limiting cross-domain effectiveness. Collectively, these findings reinforce our results on FOICE: while fine-tuning offers strong in-distribution performance, its specialization often comes at the cost of generalization, a systemic limitation shared across both audio and visual modalities.
Error patterns across detectors. An inspection of confusion matrix counts reveals divergent failure modes: baseline TCM misclassifies nearly all real samples as fake, leading to extremely high false positive rates, while base Ren et al. [AISVGeneralization] exhibits the opposite pattern with widespread false negatives. These biases underscore that detectors not only vary in performance but also in the nature of their errors, highlighting the importance of examining decision tendencies beyond aggregate metrics.
Detection under noisy and denoised conditions. Our experiments also highlight the impact of noise and audio preprocessing on detection performance. Interestingly, we observe that denoised audio is often more challenging for detectors than noisy audio. While fine-tuned AASIST and Sun et al. [VocoderArtifacts] models remain relatively stable across both conditions, detectors like TCM and Ren et al.[AISVGeneralization] show substantial degradation post-denoising. TCM’s fine-tuned variant, for example, sees accuracy drop below 90% and EER spike to 0.324 on denoised audio. This may suggest that certain detectors rely on fine-grained noise artifacts or phase distortions that are inadvertently removed by denoising algorithms, masking cues needed for detection. Future work should explore whether more adaptive or perceptually-aligned denoising techniques can retain deepfake-relevant artifacts without suppressing critical detection signals.
@@
To address these gaps, we propose a focused research agenda. First, researchers should expand generator coverage by evaluating FOICE variants integrated with alternative vocoders and video-conditioned generators to measure how vocoder choice and multi-modal conditioning affect both attack success and detector transfer. Second, studies must examine robustness under realistic channels, testing detectors across a matrix of codecs, microphone types, playback devices, and network transport conditions (including compression and packet loss) to better approximate operational environments. Finally, we encourage focused mitigation research on multi-generator / discrepancy training (cf.  [yang2025d]), domain-invariant representation learning, class-imbalance–aware fine-tuning, and perceptually-aware denoising that preserves forensic cues. These directions directly address the limitations above and point toward defensible, deployment-oriented research priorities.